OpenStudy (anonymous):

What is the value of 8!? A. 8 B. 64 C. 512 D. 40,320

OpenStudy (paki):

do you know what is this factorial...?

OpenStudy (anonymous):

8 = 8 m8.

OpenStudy (anonymous):

it A right I need somone t check my answer

OpenStudy (anonymous):

What is the value of 7P3 ? A. 21 B. 35 C. 210 D. 840

OpenStudy (anonymous):

What's there to check? 8 = 8 8^2 = 64 8^3 = 512 8^4 = 4096 ...

OpenStudy (anonymous):

What is the value of 7P3 ? A. 21 B. 35 C. 210 D. 840

OpenStudy (callmekiki):

8! = 40,320 8*7*6*5*4*3*2*1

OpenStudy (callmekiki):

7P3 = 7*6*5 Which is equal to 210

OpenStudy (callmekiki):

@clirentmaster999

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@_Squish the value of 8! isn't 8

OpenStudy (anonymous):

Ah, I see the question now. You meant "8!" and not 8!? A good way to define the two (for future reference) would be to put a space between ! and ? . Whenever you're given a number with "!," for example 8!, it means 8 x 7 x 6 x 5 x 4 x 3 x 2 x 1. The same thing goes for any number at that. 10 ! = 10 x 9 x 8 x 7... x 2 x 1.

OpenStudy (anonymous):

speak english

OpenStudy (anonymous):

What's there to expand on? Any number with "!" means that you're going to multiply from that digit to 1. 1 ! = 1 2 ! = 2 x 1 3 ! = 3 x 2 x 1 4 ! = 4 x 3 x 2 x 1 5 ! = 5 x 4 x 3 x 2 x 1 6 ! = 6 x 5 x 4 x 3 x 2 x 1 7 ! = 7 x 6 x 5 x 4 x 3 x 2 x 1 8 ! = 8 x 7 x 6 x 5 x 4 x 3 x 2 x 1 You see the pattern here, right?